# Minimal model program for log canonical pairs on complex analytic spaces

### v1

@@ 1. Introduction
Let us quickly explain these properties. The property (P2) corresponds to the affine property of the base scheme in the algebraic case. The property (P3) says that $W$ has a fundamental system of Stein open neighborhood. This property is used to shrink $Y$ around $W$ in each step of the minimal model program (MMP, for short) keeping (P2). The property (P4) is used to suppose the Noetherian property of “germ of $\mathcal{O}_{Y}$ around $W$” as in [F22a, Theorem 2.10] and the finiteness of the relative Picard rank of $\pi$ over $W$.
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,B+A)$ be an lc pair, where $B$ is an effective $\mathbb{R}$-divisor on $X$ and $A$ is an effective $\pi$-ample $\mathbb{R}$-divisor on $X$. Let $f\colon(\tilde{X},\tilde{B})\to(X,B)$ be a dlt blow-up of $(X,B)$, and we put $\tilde{\Gamma}=\tilde{B}+f^{*}A$. Let $\tilde{H}$ be a $(\pi\circ f)$-ample $\mathbb{R}$-divisor on $\tilde{X}$ such that $(\tilde{X},\tilde{\Gamma}+\tilde{H})$ is lc and $K_{\tilde{X}}+\tilde{\Gamma}+\tilde{H}$ is nef over $W$. Then there exists a sequence of steps of a $(K_{\tilde{X}}+\tilde{\Gamma})$-MMP over $Y$ around $W$ with scaling of $\tilde{H}$
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,\Delta)$ be an lc pair. Let $A$ be an effective $\mathbb{R}$-divisor on $X$ such that $(X,\Delta+A)$ is an lc pair and $K_{X}+\Delta+A$ is nef over $W$. Then no infinite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of $A$
@@
  • If we define $\lambda_{i}={\rm inf}\set{\mu\in\mathbb{R}_{\geq 0}\!}{\!K_{X_{i}}+\Delta_{i}+\mu A_{i}\text{\rm\, is nef over }W}$, where $A_{i}$ is the birational transform of $A$ on $X_{i}$, then ${\rm lim}_{i\to\infty}\lambda_{i}=0$, and
@@
For Theorem 1.3, we will establish fundamental results of the asymptotic vanishing order and the Nakayama–Zariski decomposition in [Na04] (see Section 4), techniques of MMP as in [F07] and [B12] (see Section 3), and we will apply the equi-dimensinoal reduction (cf. [AK00]) in the analytic setting established in [EH24].
@@ 2.1. Complex analytic space
Let $X$ be an analytic variety and let $\mathcal{I}$ be a coherent sheaf on $X$. Then there exists a projective bimeromorphism $f\colon X^{\prime}\to X$ from a non-singular analytic variety $X^{\prime}$ such that $\mathcal{I}\cdot\mathcal{O}_{X^{\prime}}$ is an invertible sheaf on $X^{\prime}$ and there exists an effective $f$-exceptional Cartier divisor $E^{\prime}$ on $X^{\prime}$ such that ${\rm Supp}\,E^{\prime}={\rm Ex}(f)$ and $-E^{\prime}$ is $f$-ample.
@@
[…]
@@ 2.4. Negativity lemma
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties. Let $D=E-F$ be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and there is an analytically meagre subset $S\subset{\rm Supp}\,E$ such that for any curve $C\subset{\rm Supp}\,E$ contained in a fiber of $\pi$, if $C\not\subset S$ then $(D\cdot C)\geq 0$. Then $-D\geq 0$.
@@
Then $-D+t\pi^{*}y\geq 0$, and there exist a component $P$ of $E$ and a component $Q$ of $\pi^{*}y$ such that $P\cap Q\neq\emptyset$, ${\rm coeff}_{P}(-D+t\pi^{*}y)=0$, and ${\rm coeff}_{Q}(-D+t\pi^{*}y)>0$. By cutting out $P$ with sufficiently general hyperplane sections, we can find a curve $C\subset P$ such that
@@
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties and $W\subset Y$ a subset. Let $D=E-F$ be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and there exist an $\mathbb{R}$-Cartier divisor $A$ on $X$ and a sequence of non-negative real numbers $\{\epsilon_{i}\}_{i\geq 0}$ such that ${\rm lim}_{i\to\infty}\epsilon_{i}=0$ and every $D+\epsilon_{i}A$ is movable over an open neighborhood of $W$. Then there exists an open subset $U\subset Y$ containing $W$ such that $-D|_{\pi^{-1}(U)}\geq 0$.
@@
For any $y\in W$, if there is an open subset $U_{y}\subset Y$ containing $y$ such that $-D|_{\pi^{-1}(U_{y})}\geq 0$, then $U:=\bigcup_{y\in W}U_{y}$ is the desired open subset. Thus, we may assume that $W$ is a point. By shrinking $Y$ around $W$, we may assume that $Y$ is Stein and there is a Zariski open subset $Y^{\prime}\subset Y$ such that the image of all components of $D^{\prime}:=D|_{\pi^{-1}(Y^{\prime})}$ contain $W$.
By definition of movable divisors, there is an open subset $U_{i}\subset Y^{\prime}$ containing $W$ such that any irreducible component of ${\boldsymbol{\rm Bs}}|(D+\epsilon_{i}A)|_{\pi^{-1}(U_{i})}/U_{i}|_{\mathbb{R}}$ has codimension $\geq 2$. Since $Y$ is Stein, by Proposition 2.9 and the fact that the image of all components of $D^{\prime}$ contain $W$, any component of $E^{\prime}:=E|_{\pi^{-1}(Y^{\prime})}$ is not contained in ${\boldsymbol{\rm Bs}}|D+\epsilon_{i}A|_{\mathbb{R}}$. Therefore, we can find an analytically meagre subset $S\subset{\rm Supp}\,E^{\prime}$ such that any point $x\in{\rm Supp}\,E^{\prime}\setminus S$ is not contained in any ${\boldsymbol{\rm Bs}}|D+\epsilon_{i}A|_{\mathbb{R}}$. From this, for any curve $C\subset{\rm Supp}\,E^{\prime}$ contained in a fiber of $\pi$, if $C\not\subset S$ then $(D\cdot C)={\rm lim}_{i\to\infty}(D+\epsilon_{i}A)\cdot C\geq 0$. By Lemma 2.14, we have $-D|_{\pi^{-1}(Y^{\prime})}\geq 0$. Thus, Corollary 2.15 holds. ∎
@@
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties and $W\subset Y$ a subset. Let $D=E-F$ be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and $D$ is nef over $W$. Then, there is an open subset $U\supset W$ of $Y$ such that $-D|_{\pi^{-1}(U)}\geq 0$. In particular, for any bimeromorphism $\pi\colon X\to Y$ of normal analytic varieties and any $\mathbb{R}$-Cartier divisor $D$ on $X$ such that $D$ is nef over $W$ and $-\pi_{*}D\geq 0$, there exists an open subset $U\subset Y$ containing $W$ such that such that $-D|_{\pi^{-1}(U)}\geq 0$.
@@
For any $y\in W$, if there is an open subset $U_{y}\subset Y$ containing $y$ such that $-D|_{\pi^{-1}(U_{y})}\geq 0$, then $U:=\bigcup_{y\in W}U_{y}$ is the desired open subset. Thus, we may assume that $W$ is a point. Let $A$ be a $\pi$-ample Cartier divisor on $X$. By [F22a, Lemma 4.10], for any $\epsilon\in\mathbb{R}_{>0}$, there is a Stein open subset $U_{\epsilon}\subset Y$ containing $W$ such that $D+\epsilon A$ is ample over $U_{\epsilon}$. In particular, ${\boldsymbol{\rm Bs}}|(D+\epsilon A)|_{\pi^{-1}(U_{\epsilon})}/U_{\epsilon}|_{\mathbb{R}}$ is empty. Corollary 2.16 follows from this fact and Corollary 2.15. ∎
@@ 2.5. Singularities of pairs
[…]
@@ 2.7. Bimeromorphic map
Let $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y$ be projective morphisms from normal analytic varieties to an analytic space $Y$, and let $\phi\colon X\dashrightarrow X^{\prime}$ be a bimeromorphic contraction over $Y$. Let $D$ be an $\mathbb{R}$-Cartier divisor on $X$ such that $-D$ is ample over $Y$. Suppose that $\phi_{*}D$ is $\mathbb{R}$-Cartier and ample over $Y$. Let $g\colon W\to X$ and $g^{\prime}\colon W\to X^{\prime}$ be the normalization of the graph of $\phi$. We put
@@
Let $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y$ be projective morphisms from normal analytic varieties to an analytic space $Y$, and let $\phi\colon X\dashrightarrow X^{\prime}$ be a bimeromorphic contraction over $Y$. Let $(X,\Delta)$ and $(X^{\prime},\Delta^{\prime})$ be dlt pairs such that $\Delta^{\prime}=\phi_{*}\Delta$. Suppose that $-(K_{X}+\Delta)$ and $K_{X^{\prime}}+\Delta^{\prime}$ are ample over $Y$. Then the following hold.
@@ 3.1. Models
  • $\rho(X^{\prime}/Y;W)-\rho(Z/Y;W)=1$ and $-(K_{X^{\prime}}+\Delta^{\prime})$ is ample over $Z$, and
@@
If $W=\emptyset$, then we formally define $(X,\Delta)$ itself to be a log minimal model (and thus a good minimal model) of $(X,\Delta)$ over $Y$ around $W$. Let $(\widetilde{X},\widetilde{\Delta})$ be a disjoint union $\coprod_{\lambda\in\Lambda}(\widetilde{X}_{\lambda},\widetilde{\Delta}_{\lambda})$ of lc pairs, $\widetilde{\pi}\colon\widetilde{X}\to Y$ a projective morphism to an analytic space $Y$, and $W\subset Y$ a subset. Then a log birational model of $(\widetilde{X},\widetilde{\Delta})$ over $Y$ (resp. a log minimal model of $(\widetilde{X},\widetilde{\Delta})$ over $Y$ around $W$) is a disjoint union of lc pairs $\coprod_{\lambda\in\Lambda}(\widetilde{X}^{\prime}_{\lambda},\widetilde{\Delta}^{\prime}_{\lambda})$ such that $(\widetilde{X}^{\prime}_{\lambda},\widetilde{\Delta}^{\prime}_{\lambda})$ is a log birational model of $(\widetilde{X}_{\lambda},\widetilde{\Delta}_{\lambda})$ over $Y$ (resp. a log minimal model of $(\widetilde{X}_{\lambda},\widetilde{\Delta}_{\lambda})$ over $Y$ around $W$) for all $\lambda\in\Lambda$.
@@
for $i=1$ and $2$, where $P$ runs over prime divisors on $X^{\prime}$. By the definition of weak lc models, $E_{i}$ is effective. When a prime divisor $P$ on $X^{\prime}$ is not exceptional over $X$ or $X_{i}$, obviously ${\rm coeff}_{P}(E_{i})=0$. When $P$ is exceptional over $X$ but not exceptional over $X_{i}$, then $a(P,X_{i},\Delta_{i})=-1$. Then $-1=a(P,X_{i},\Delta_{i})\geq a(P,X,\Delta)\geq-1$ since $(X,\Delta)$ is lc. Therefore, we have ${\rm coeff}_{P}(E_{i})=0$. From these facts, $E_{i}$ is $f_{i}$-exceptional.
@@
The proof is the same as the algebraic case. Let $(\widetilde{X},\widetilde{\Delta})$ be an lc pair. By the negativity lemma (Corollary 2.16), we can easily check that $(\widetilde{X},\widetilde{\Delta})$ is a weak lc model of $(X,\Delta)$ over $Y$ around $W$ if and only of $(\widetilde{X},\widetilde{\Delta})$ is a weak lc model of $(X^{\prime},\Delta^{\prime})$ over $Y$ around $W$. We omit the details. ∎

### v2

@@ 1. Introduction
Let us quickly explain these properties. The property (P2) corresponds to the affine property of the base scheme in the algebraic case. The property (P3) says that $W$ has a fundamental system of Stein open neighborhood. This property is used to shrink $Y$ around $W$ in each step of the minimal model program (MMP, for short) keeping (P2). The property (P4) is used to suppose the Noetherian property of “germ of $\mathcal{O}_{Y}$ around WW” as in [F22a, Theorem 2.10] and the finiteness of the relative Picard rank of $\pi$ over $W$.
@@
[…]
@@
Theorem 1.2 is new and nontrivial even in the klt case because we cannot directly apply the argument by Lai [Lai11] to the $\mathbb{R}$-boundary divisor case, and we need the equi-dimensinoal reduction (cf. [AK00]) in the analytic setting established in [EH24].
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to a Stein space $Y$, and let $W\subset Y$ be a compact subset such that $\pi$ and $W$ satisfy (P). Let $(X,\Delta)$ be an lc pair. Let $A$ be an effective $\mathbb{R}$-divisor on $X$ such that $(X,\Delta+A)$ is an lc pair and $K_{X}+\Delta+A$ is nef over $W$. Then no infinite sequence of steps of a $(K_{X}+\Delta)$-MMP over $Y$ around $W$ with scaling of AA
@@
[…]
@@
For Theorem 1.3, we will establish fundamental results of the asymptotic vanishing order and the Nakayama–Zariski decomposition in [Na04] (see Section 4), techniques of MMP as in [F07] and [B12] (see Section 3), and we will apply the equi-dimensinoal reduction (cf. [AK00]) in the analytic setting established in [EH24]. We also note that some important results on the log canonical abundance in the algebraic case are necessary to prove the main results of this paper. For example, to prove Lemma 5.4, which is crucial to the proof of Theorem 1.3, we need the abundance theorem for lc pairs of numerical dimension zero [G13] by Gongyo. Moreover, we need [HH20, Theorem 1.5] by the second author and Hu for the reduction of Theorem 1.1 to Theorem 1.3.
@@ 2.1. Complex analytic space
Let $X$ be an analytic variety and let $\mathcal{I}$ be a coherent sheaf on $X$. Then there exists a projective bimeromorphism $f\colon X^{\prime}\to X$ from a non-singular analytic variety $X^{\prime}$ such that $\mathcal{I}\cdot\mathcal{O}_{X^{\prime}}$ is an invertible sheaf on $X^{\prime}$ and there exists an effective $f$-exceptional Cartier divisor $E^{\prime}$ on $X^{\prime}$ such that ${\rm Supp}\,E^{\prime}={\rm Ex}(f)$ and $E^{\prime}$ is $f$-ample.
@@
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to an analytic space $Y$, and let $W\subset Y$ be a compact subset such that $W\cap Z$ has only finitely many connected components for any analytic subset $Z$ which is defined over an open neighborhood of $W$ (see also the condition $({\rm P}4)$ in [F22a]). Then there exists an open subset $Y^{\prime}\subset Y$ containing $W$ such that if $\pi^{-1}(Y^{\prime})=\coprod_{\lambda\in\Lambda}X^{\prime}_{\lambda}$ is the decomposition of $\pi^{-1}(Y^{\prime})$ into connected components $X^{\prime}_{\lambda}$, then there are only finitely many connected components, which we denote by $X^{\prime}_{1},\cdots,\,X^{\prime}_{l}$, such that their images on $Y$ by $\pi$ intersect $W$. Furthermore, these $X^{\prime}_{1},\cdots,\,X^{\prime}_{l}$ satisfy the following properties.
@@ 2.4. Negativity lemma
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties. Let D=EFD=E-F be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and there is an analytically meagre subset $S\subset{\rm Supp}\,E$ such that for any curve $C\subset{\rm Supp}\,E$ contained in a fiber of $\pi$, if $C\not\subset S$ then $(D\cdot C)\geq 0$. Then $D\geq 0$.
@@
Then $D+t\pi^{*}y\geq 0$, and there exist a component $P$ of $E$ and a component $Q$ of $\pi^{*}y$ such that $P\cap Q\neq\emptyset$, ${\rm coeff}_{P}(-D+t\pi^{*}y)=0$, and ${\rm coeff}_{Q}(-D+t\pi^{*}y)>0$. By cutting out $P$ with sufficiently general hyperplane sections, we can find a curve $C\subset P$ such that
@@
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties and $W\subset Y$ a subset. Let D=EFD=E-F be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and there exist an $\mathbb{R}$-Cartier divisor $A$ on $X$ and a sequence of non-negative real numbers $\{\epsilon_{i}\}_{i\geq 0}$ such that ${\rm lim}_{i\to\infty}\epsilon_{i}=0$ and every $D+\epsilon_{i}A$ is movable over an open neighborhood of $W$. Then there exists an open subset $U\subset Y$ containing $W$ such that $D|_{\pi^{-1}(U)}\geq 0$.
@@
For any $y\in W$, if there is an open subset $U_{y}\subset Y$ containing $y$ such that $D|_{\pi^{-1}(U_{y})}\geq 0$, then $U:=\bigcup_{y\in W}U_{y}$ is the desired open subset. Thus, we may assume that $W$ is a point. By shrinking $Y$ around $W$, we may assume that $Y$ is Stein and there is a Zariski open subset $Y^{\prime}\subset Y$ such that the image of all components of $D^{\prime}:=D|_{\pi^{-1}(Y^{\prime})}$ contain $W$.
By definition of movable divisors, there is an open subset $U_{i}\subset Y^{\prime}$ containing $W$ such that any irreducible component of ${\boldsymbol{\rm Bs}}|(D+\epsilon_{i}A)|_{\pi^{-1}(U_{i})}/U_{i}|_{\mathbb{R}}$ has codimension $\geq 2$. Since $Y$ is Stein, by Proposition 2.9 and the fact that the image of all components of $D^{\prime}$ contain $W$, any component of $E^{\prime}:=E|_{\pi^{-1}(Y^{\prime})}$ is not contained in ${\boldsymbol{\rm Bs}}|D+\epsilon_{i}A|_{\mathbb{R}}$. Therefore, we can find an analytically meagre subset $S\subset{\rm Supp}\,E^{\prime}$ such that any point $x\in{\rm Supp}\,E^{\prime}\setminus S$ is not contained in any ${\boldsymbol{\rm Bs}}|D+\epsilon_{i}A|_{\mathbb{R}}$. From this, for any curve $C\subset{\rm Supp}\,E^{\prime}$ contained in a fiber of $\pi$, if $C\not\subset S$ then $(D\cdot C)={\rm lim}_{i\to\infty}(D+\epsilon_{i}A)\cdot C\geq 0$. By Lemma 2.14, we have $D|_{\pi^{-1}(Y^{\prime})}\geq 0$. Thus, Corollary 2.15 holds. ∎
@@
Let $\pi\colon X\to Y$ be a contraction of normal analytic varieties and $W\subset Y$ a subset. Let D=EFD=E-F be an $\mathbb{R}$-Cartier divisor on $X$, where $E$ and $F$ are effective $\mathbb{R}$-divisors having no common components. Suppose that $E$ is very exceptional over $Y$ and $D$ is nef over $W$. Then, there is an open subset $U\supset W$ of $Y$ such that $D|_{\pi^{-1}(U)}\geq 0$. In particular, for any bimeromorphism $\pi\colon X\to Y$ of normal analytic varieties and any $\mathbb{R}$-Cartier divisor $D$ on $X$ such that $D$ is nef over $W$ and $\pi_{*}D\geq 0$, there exists an open subset $U\subset Y$ containing $W$ such that such that $D|_{\pi^{-1}(U)}\geq 0$.
@@
For any $y\in W$, if there is an open subset $U_{y}\subset Y$ containing $y$ such that $D|_{\pi^{-1}(U_{y})}\geq 0$, then $U:=\bigcup_{y\in W}U_{y}$ is the desired open subset. Thus, we may assume that $W$ is a point. Let $A$ be a $\pi$-ample Cartier divisor on $X$. By [F22a, Lemma 4.10], for any $\epsilon\in\mathbb{R}_{>0}$, there is a Stein open subset $U_{\epsilon}\subset Y$ containing $W$ such that $D+\epsilon A$ is ample over $U_{\epsilon}$. In particular, ${\boldsymbol{\rm Bs}}|(D+\epsilon A)|_{\pi^{-1}(U_{\epsilon})}/U_{\epsilon}|_{\mathbb{R}}$ is empty. Corollary 2.16 follows from this fact and Corollary 2.15. ∎
@@ 2.5. Singularities of pairs
Let $\pi\colon X\to Y$ be a projective morphism from a normal analytic variety $X$ to an analytic space $Y$, and let $W\subset Y$ be a compact subset such that $W\cap Z$ has only finitely many connected components for any analytic subset $Z$ which is defined over an open neighborhood of $W$ (see also the condition $({\rm P}4)$ in [F22a]). Let $(X,\Delta)$ be a dlt pair. Then there exists an open subset $Y^{\prime}\subset Y$ containing $W$ such that every connected component $X^{\prime}$ of $\pi^{-1}(Y^{\prime})$ satisfying the following conditions.
@@ 2.7. Bimeromorphic map
Let $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y$ be projective morphisms from normal analytic varieties to an analytic space $Y$, and let $\phi\colon X\dashrightarrow X^{\prime}$ be a bimeromorphic contraction over $Y$. Let $D$ be an $\mathbb{R}$-Cartier divisor on $X$ such that D-D is ample over $Y$. Suppose that $\phi_{*}D$ is $\mathbb{R}$-Cartier and ample over $Y$. Let $g\colon W\to X$ and $g^{\prime}\colon W\to X^{\prime}$ be the normalization of the graph of $\phi$. We put
@@
Let $f\colon X\to Y$ and $f^{\prime}\colon X^{\prime}\to Y$ be projective morphisms from normal analytic varieties to an analytic space $Y$, and let $\phi\colon X\dashrightarrow X^{\prime}$ be a bimeromorphic contraction over $Y$. Let $(X,\Delta)$ and $(X^{\prime},\Delta^{\prime})$ be dlt pairs such that $\Delta^{\prime}=\phi_{*}\Delta$. Suppose that $(K_{X}+\Delta)$ and $K_{X^{\prime}}+\Delta^{\prime}$ are ample over $Y$. Then the following hold.
@@ 3.1. Models
  • $\rho(X^{\prime}/Y;W)-\rho(Z/Y;W)=1$ and $(K_{X^{\prime}}+\Delta^{\prime})$ is ample over $Z$, and
@@
[…]
@@
for $i=1$ and $2$, where $P$ runs over prime divisors on $X^{\prime}$. By the definition of weak lc models, $E_{i}$ is effective. When a prime divisor $P$ on $X^{\prime}$ is not exceptional over $X$ or $X_{i}$, obviously ${\rm coeff}_{P}(E_{i})=0$. When $P$ is exceptional over $X$ but not exceptional over $X_{i}$, then $a(P,X_{i},\Delta_{i})=-1$. Then $1=a(P,X_{i},\Delta_{i})\geq a(P,X,\Delta)\geq-1$ since $(X,\Delta)$ is lc. Therefore, we have ${\rm coeff}_{P}(E_{i})=0$. From these facts, $E_{i}$ is $f_{i}$-exceptional.
@@
[…]